How to find the total monthly sum?
Names: Amélie and Félix Rodrigue, 24 years old.
Current situation: Amélie and Félix live in an apartment. They plan to buy a house in five and a half years. They would like to be able to increase their current down payment to the maximum.
They would like to accumulate $ 45,000 to add to the down payment.
Gross family income: $ 152,128
Salary increases confirmed: 2% per year for the next five years.
Down payment accumulated to date: $ 30,000
Children: the family has no children yet. They plan to have two children after buying the house.
Wishes: Amélie would like to live in a neighborhood less than 10 years old.
Possible investments: 6-month certificates at 3% per year compounded monthly.
1
Expert's answer
2020-11-08T18:49:43-0500

Gross family income is the sum of annual gross pay plus interest on certificates.

P'=P*(1+r/n)nt - Compound interest formula

Let total annual sum be x.

152128 = x + P*(1+r/n)nt-P

152128 = x + 30000*(1+0.03/12)12-30000

152128 = x + 913

x = 151215

total monthly sum = total annual sum/12 = 12601

Answer: total monthly sum = 12601 $